OpenStudy (isaiah.feynman):

How to simplify the square root of 126

OpenStudy (dape):

\[\sqrt{126}=\sqrt{3^2*2*7}=3*\sqrt{14}\] Where I factored 126 into prime factors in the second step.

OpenStudy (anonymous):

\[ \sqrt{126} = \sqrt{63 \times 2} = \sqrt{9 \times 7 \times 2} = 3\sqrt{ 7 \times 2} \] \[= 3\sqrt{ 14} \]

OpenStudy (isaiah.feynman):

Thanks!